OpenStudy (anonymous):

Using the greatest possible error find the difference between the maximum and minimum areas of a garden measuring 18 ft by 36 ft.

ganeshie8 (ganeshie8):

the given measurements are rounded to nearest unit place, that means greatest possible error = \(\pm 0.5\)

ganeshie8 (ganeshie8):

calculate area wid minimum dimensions calculate area wid maximum dimensions subtracct

ganeshie8 (ganeshie8):

Area wid minimum dimensions = \(17.5 \times 35.5\)

ganeshie8 (ganeshie8):

Area wid maximum dimensions = \(18.5 \times 36.5\)

ganeshie8 (ganeshie8):

difference in areas = \(\large (18.5 \times 36.5) - (17.5 \times 35.5)\)

ganeshie8 (ganeshie8):

plugin to ur calculator

ganeshie8 (ganeshie8):

let me knw if somthing doesnt make sense :)

OpenStudy (anonymous):

totally dont get it

OpenStudy (anonymous):

ty though

ganeshie8 (ganeshie8):

which part is throwing u off ?

ganeshie8 (ganeshie8):

never heard of 'greatest possible error' before ?

OpenStudy (anonymous):

right

ganeshie8 (ganeshie8):

http://www.phschool.com/atschool/academy123/english/academy123_content/wl-book-demo/ph-170s.html go thru that short video, and u may come back and ask if u dont get something. be specific ok ? saying 'i totally dont get it' wont give me clue where exactly u have problem in understanding :/

OpenStudy (anonymous):

tnx watched video ...i understand it better/:)

ganeshie8 (ganeshie8):

good to hear :) so wil u try going thru above work.. and see if it looks okay now ha ?

OpenStudy (anonymous):

yes got 54

OpenStudy (anonymous):

ty

ganeshie8 (ganeshie8):

let me check quick..

ganeshie8 (ganeshie8):

\(54\) is correct ! good job !! http://www.wolframalpha.com/input/?i=+%2818.5++%5Ctimes++++36.5%29+-+++%2817.5++%5Ctimes++++35.5%29